OpenStudy (anonymous):

A circle has its center at the center of a square with 8-inch sides. Find the area of the square not covered by the circle. Round the answer to the nearest tenth

OpenStudy (anonymous):

Are of square: 8 x 8 = Area of circle: Pi*R^2 = 3.14*(4)^2 = Difference of area: .......